OpenStudy (anonymous):

Yara rolls a standard die and flips a coin. What is the probability that she gets an even number and heads? twenty-five percent fifty percent seventy-five percent one hundred percent

OpenStudy (anonymous):

1.) 25%

OpenStudy (anonymous):

(3/6) x (1/2) = 3/12 = 1/4= A) 25%

OpenStudy (anonymous):

did that help you? if it did please mark me as your "best answer" it would be greatly appreciated.